For an enzyme that displays Michaelis-Menten kinetics, what is the reaction velocity v (as a percentage of Vmax) ,observed at each of the following substrate concentrations. (Ex, v = xVmax, where x = an integer, fraction, or decimal number (two decimal places))

a) [S] = 0.1 Km _________

b) [S] = 2 Km _________

c) [S] = 10 Km _________
Answer #1


V = I x V mar - [S] KM + [] * Vmar :: 1 = KM+ (Si [S]

a)

S = 0.1KM KM + [S] 0.1KM = K +0.1KM 0.1KM T = (1 + 0.1 KM

0.1KM 1.1KM 0.1 1 = 1.1 Τ = 0,0909

b)

(s) = 2KM (87 KM + [S] 2KM KM + 2KM 2KM (1+2) KM

2км 3КМ т= 0,667

C)

(s) = 10KM 151 KM + [S] 10KM <= KM + 10KM 10KM (1 +10) KM

6060 = 1 II - 1 10 11KM 10KM = x